OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = 2x + 3 and the given point (-4,1). 12 Answer: y=2x

OpenStudy (anonymous):

@Princer_Jones can you please show me how to do this

OpenStudy (anonymous):

Slope of parallel lines are equal. The slope of the given line is 2 . Hence we have to find the equation of the line with slope 2 and having (-4,1) as a point . The required equation is given by y-1=2(x+4)this implies y-1=2x+8 this implies y=2x+9 @omg1help

OpenStudy (anonymous):

IN the question , what is it said about the point (-4,1)? @omg1help

OpenStudy (anonymous):

it the given point on a graph

OpenStudy (anonymous):

can you post the question again? @omg1help

OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = 2x + 3 and the given point (-4,1). 12

OpenStudy (anonymous):

where is the 12??? @omg1help

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

I can solve any problem of this course, but you need to provide me a proper question, the 12 must be plugged in somewhere . @omg1help

OpenStudy (anonymous):

no it was just the number of the next question